Grade 12Trigonometry

Tan a =1/2 ,tan b = 1/5, tan c = 1/8 then find tan( a + b + c) =

Bhanu prakash
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
Dear Bhanu
 
tan a = 1/2
tan b = 1/5
tan(a+b) = (tan a + tan b) /(1 - tan a tan b)
tan(a+b) = (1/2 + 1/5)/(1 - 1/2. 1/5)
tan(a+b) = (5+2)/(10- 1)
tan(a+b) = 7/9
tan (a+b+c) =  tan {(a+b) + c}
= { tan(a+b) + tan c } / { 1 - tan(a+b).tan c }
= {7/9 + 1/8) / (1 - 7/9.*1/8)
= (56 + 9) / (72- 7)
= 65/65
= 1